A force of 〖3.0x10〗^(-6) N exists between two charged objects whose centers are 3.0m apart. One of the objects has a charge of 〖1.5x10〗^(-7) C. Find the charge on the other object. 



1
Expert's answer
2021-04-19T09:04:55-0400

According to Coulomb's law expression:


"F=k\frac{q_1q_2}{r^2},\\\space\\
q_2=\frac{Fr^2}{kq_1}=20·10^{-9}\text{ C}."